OpenStudy (anonymous):

The adjacent sides of a parallelogram measures 14 centimeters in 20 centimeters and one angle of 57 degrees what is the area of the parallelogram

jimthompson5910 (jim_thompson5910):

 Area of parallelogram: \[\Large A = a*b*\sin(\theta)\]